The TV repair company receives TVs with various faults. Therefore, their correction time Tis a random variable. Find the probability density function of this random dudge if the probability distribution function F(t) has the following form: (0, t <0, F(t) ={1-, t>0,


1
Expert's answer
2021-01-05T18:28:04-0500

"F(t) = \begin{cases}
 0, &t<0 \\
 1-e^{-t}, &t\geq0
\end{cases}"

"f(t)=\dfrac{dF}{dt}"

"f(t) = \begin{cases}
 0, &t<0 \\
 e^{-t}, &t\geq0
\end{cases}"

"\displaystyle\int_{-\infin}^{\infin}f(t)dt=\displaystyle\int_{0}^{\infin}e^{-t}dt"

"=\lim\limits_{A\to \infin}[-e^{-t}]\begin{matrix}
 A \\
 0
\end{matrix}=-0+1=1"

"f(t) = \begin{cases}
 0, &t<0 \\
 e^{-t}, &t\geq0
\end{cases}"